# High Performance Space Debris Tracking in Complex Skylight Backgrounds with a Large-Scale Dataset

### v3

@@ Abstract
With the rapid development of space exploration, space debris has attracted more attention due to its potential extreme threat, leading to the need for real-time and accurate debris tracking. However, existing methods are mainly based on traditional signal processing, which cannot effectively process the complex background and dense space debris. In this paper, we propose a deep learning-based Space Debris Tracking Network (SDT-Net) to achieve highly accurate debris tracking. SDT-Net effectively represents the feature of debris, enhancing the efficiency and stability of end-to-end model learning. To train and evaluate this model effectively, we also produce a large-scale dataset Space Debris Tracking Dataset (SDTD) by a novel observation-based data simulation scheme. SDTD contains 18,040 video sequences with a total of 62,562 frames and covers 250,000 synthetic space debris. Extensive experiments validate the effectiveness of our model and the challenging of our dataset. Furthermore, we test our model on real data from the Antarctic Station, achieving a MOTA score of 70.6%, which demonstrates its strong transferability to real-world scenarios. Our dataset and code will be released soon.
@@ 1 Introduction
During the past decade, global space activities have exponentially increased, which has raised the potential risk of collisions between debris and spacecraft in near-Earth space [1, 2].
This has led to significant economic losses, drawing increasing attention to the issue of space debris. The Space Debris Tracking (SDT) is to detect and predict the trajectories of space debris, reducing the collision risks and promoting space industry development. This makes monitoring debris extremely urgent and critical for ensuring the safety of space operations.
SDT is an essential task that includes detecting and predicting the trajectories of space debris. It is a computer vision problem, specifically an object tracking task in astronomical videos, where the goal is to identify and follow space debris across continuous frames for a video. However, traditional methods typically focus on object detection rather than tracking. These approaches often rely on techniques such as template-matching methods [3, 4] and morphological operator methods [5, 6]. While these methods are capable of detecting debris in individual images, they struggle to associate objects across consecutive frames, making them less suitable for space debris monitoring.
@@
To address the challenge of limited data in astronomy, we follow the idea of observation-based data simulation in the astronomical field [11, 12, 13]. The observation-based data simulation is widely proven to lead to realistic data and reliable scientific conclusions, providing a chance for us to design a method for data simulating space debris. Due to its success, we construct a Space Debris Tracking Dataset (SDTD), which is the first publicly available dataset designed for space debris tracking. Tab 1 highlights the key differences of SDTD compared to existing datasets. Specifically, as shown in Fig 1, SDTD consists of 18,040 synthetic videos, totaling 62,562 frames and 25,000 annotated debris instances.
In order to apply AI methods to the SDT task, we propose a simple yet effective deep learning method, namely the Space Debris Tracking Network (SDT-Net). We introduce a Region-of-Interest Feature Enhancement (RoIFE) module to highlight debris features, aiding in more accurate detection. A detection module is then used to localize the debris by predicting their positions in the image. After localization, a tracking module is applied to perform cross-frame data association, generating tracking trajectories for each debris.
@@
[11] Real 83 images 7 video $\checkmark$ $\times$ $\times$ One One
@@
[6] Real 200 images $\times$ $\checkmark$ $\times$ $\times$ One One
@@ 3 Benchmark Dataset
In this paper, we utilize astronomical images collected by the Zwicky Transient Facility (ZTF) [29] to simulate space debris.
We propose a simulation tool to generate space debris, which provides sufficient data for constructing the SDTD dataset. Under long-exposure conditions in telescopic observations, debris often appears as a line source in the images. So, our simulation target is to generate realistic long-exposure line sources. We develop a pipeline that includes data processing, line source debris initialization, point diffusion function process, and video generation. Fig 3 and the following paragraphs provide a detailed description of this process.
@@ 3.1.2 Data Pre-Processing
where $z_{1}$ and $z_{2}$ represent the lower and upper bounds of the pixel intensity range, respectively. The hyperparameter $k$ is assigned a value of 2.5. Then, we perform a linear stretch on the grayscale values of the image, mapping the range $[z_{1},z_{2}]$ to $[0,255]$:
@@
$I_{final}(x,y)=min((max(I_{scaled}(x,y),0),255).$ (3)
@@ 3.2.2 Post Processing for Realistic
Since the real line source debris conforms to the Point Diffusion Function (PSF) characteristics [31], we perform point diffusion function processing on the preliminary line source debris. Among them, PSF is expressed as follows:
@@
Figure 3: Overview of the space debris simulation pipeline. Step 1: Data Collection. Step 2: ZTF Raw Data Filtering. Step 3: Simulation Tool for Space Debris. Various parameters, including random initial points, angle change, length change, point spread function (PSF), fracture interference, and displacement, are applied to simulate space debris. The process generates synthetic video sequences and shows the ground truth for reference.
@@ 3.3.1 Dataset Statistics
After the simulation, we obtain 18,040 videos with a total of 37,829 frames. As shown in Fig. 2(a), to evaluate model performance in complex environments, we select 1,000 images from the 16,040 dataset to create two test sets: the debris test set and the dense debris test set. These two test sets allow for a comprehensive evaluation of model robustness under different conditions, including dense debris.
@@ 4.1 Overview
The entire pipeline of our Space Debris Tracking Network (SDT-Net) method is shown in Fig. 4, which is a CenterTrack-style [21] tracking method. Due to the debris exposure as linear, we treat the debris tracking as tracking for linear segments. For each object, we follow the tracking-by-detection idea. Our model detects two endpoints of each line and predicts velocity for cross-frame object association to generate trajectory.
@@ 4.3 Region-of-Interest Feature Enhancement
where $W,H$ are image dimensions, $M_{ij}$ is the ground truth, and $\hat{M}_{ij}$ is the predicted probability.
The obtained $F_{en}$ is then passed through two modules: a Debris Detection module and a Debris Offset module.
@@ 5 Experiments
Table 2: Comparison of trackers on Debris Test and Dense Debris Test. The table compares the performance of various tracking algorithms in two distinct scenarios.
Tracker Debris Dense Debris
@@ 5.1 Experimental Setup
All experiments are performed using the mmdetection framework [39] with Pytorch [40]. The SDT-Net model is trained on 8 NVIDIA 4090 GPUs. The entire model is trained for 60 epochs, and the initial learning rate is set to 3e-3, which is decayed by 0.1 at the 20 epoch. The association radius $r$ to 200. Additionally, We adopt a downsampling strategy to generate paired images for training. Specifically, we first crop image patches of size 1524$\times$1524 from the original high-resolution images as ground truth. We employ
Figure 5: Visualization of sample tracking results of SDTD on the test set. The green box represents gt and the red box represents the predicted box.
the resolution of the input images is set to 1524$\times$1524 and the batch size is set to 2. Besides, the value of $\lambda_{\text{seg}}$ and $\lambda_{\text{emb}}$ are set to 1.0, $\lambda_{\text{hm}}$ is set to 10, and $\lambda_{\text{off}}$ is set to 0.1,
@@ 5.2 Evaluation Metrics
For evaluating the performance, we use standard MOT [41] evaluation metrics. Specifically, we adopt commonly used metrics in the field, including Multi-Object Tracking Accuracy (MOTA), identity F1 score (IDF1), Identity Switches (IDs), and Higher Order Tracking Accuracy (HOTA) [42]. These metrics provide a comprehensive assessment of tracking performance.
@@ 5.4 Ablation Study
LSE RoI-FE Offset Module Debris Test Dense Debris Test
@@
Table 4: Evaluation of Antarctic station data on SDT-Net.
[…]
CenterTrack 64.7 58.3 31 1895
OCSORT 69.3 66.8 31 1895
[…]
Figure 6: Visualization of sample tracking results of real observations.
@@ 5.5 Qualitative Results
To further validate our approach, we invited an astronomy expert to annotate real observational data. From the available real-world dataset, we randomly selected five video sequences comprising a total of 333 frames for this user study. As shown in Table 4, our method (SDT-Net) achieves a MOTA score exceeding 70%, demonstrating strong real-world performance.
Fig 6 further illustrates the effectiveness of our approach. The left side of the figure presents raw debris observations, while the right side visualizes the predicted tracking results. Notably, the right side does not represent ground truth but rather our model’s predictions. For clarity, we display each tracked object using a line connecting its two key points instead of conventional bounding boxes. In practical applications, these tracked trajectories are overlaid on blank background images to facilitate interpretation. Each color in the visualization corresponds to a distinct debris ID, reflecting the model’s ability to distinguish and track multiple objects.

### v4

@@ Abstract
With the rapid development of space exploration, space debris has attracted more attention due to its potential extreme threat, leading to the need for real-time and accurate debris tracking. However, existing methods are mainly based on traditional signal processing, which cannot effectively process the complex background and dense space debris. In this paper, we propose a deep learning-based Space Debris Tracking Network (SDT-Net) to achieve highly accurate debris tracking. SDT-Net effectively represents the feature of debris, enhancing the efficiency and stability of end-to-end model learning. To train and evaluate this model effectively, we also produce a large-scale dataset Space Debris Tracking Dataset (SDTD) by a novel observation-based data simulation scheme. SDTD contains 18,040 video sequences with a total of 62,562 frames and covers 250,000 synthetic space debris. Extensive experiments validate the effectiveness of our model and the challenging of our dataset. Furthermore, we test our model on real data from the Antarctic Station, achieving a MOTA score of 73.2%, which demonstrates its strong transferability to real-world scenarios. Our dataset and code will be released soon.
@@ 1 Introduction
During the past decade, global space activities have exponentially increased, which has raised the potential risk of collisions between debris and spacecraft in near-Earth space [1, 2]. This has led to significant economic losses, drawing increasing attention to the issue of space debris. The Space Debris Tracking (SDT) is to detect and predict the trajectories of space debris, reducing the collision risks and promoting space industry development. This makes monitoring debris extremely urgent and critical for ensuring the safety of space operations.
SDT is an essential task that includes detecting and predicting the trajectories of space debris. It is a computer vision problem, specifically an object tracking task in astronomical videos, where the goal is to identify and follow space debris across continuous frames for a video. However, traditional methods typically focus on object debris detection rather than tracking. These approaches often rely on techniques such as template-matching methods [3, 4] and morphological operator methods [5, 6]. While these methods are capable of detecting debris in individual images, they struggle to associate objects across consecutive frames, making them less suitable for space debris monitoring.
@@
To address the challenge of limited data in astronomy, we follow the idea of observation-based data simulation in the astronomical field [11, 12, 13]. The observation-based data simulation is widely proven to lead to realistic data and reliable scientific conclusions, providing a chance for us to design a method for data simulating space debris. Due to its success, we construct a Space Debris Tracking Dataset (SDTD), which is the first publicly available dataset designed for space debris tracking. Fig 1 shows the formation principle, type, examples, and corresponding SDT tasks of space debris. Specifically, Tab 1 highlights the key differences of SDTD compared to existing datasets. SDTD consists of 18,040 synthetic videos, totaling 62,562 frames and 25,000 annotated debris instances.
In order to apply AI methods to the SDT task, we propose a simple yet effective deep learning method, namely the Space Debris Tracking Network (SDT-Net). We introduce a Region-of-Interest Feature Enhancement (RoIFE) module to highlight debris features, aiding in more accurate detection. A detection module is then used to localize the debris by predicting their positions in the image. After localization, a debris offset module is applied to perform cross-frame data association, generating tracking trajectories for each debris.
@@
Lin et al. [11] Real 83 images 7 video $\checkmark$ $\times$ $\times$ One One
@@
Jiang et al. [6] Real 200 images $\times$ $\checkmark$ $\times$ $\times$ One One
@@ 3 Benchmark Dataset
In this paper, we utilize astronomical images collected by the Zwicky Transient Facility (ZTF) [29] to simulate space debris. We propose a simulation tool to generate space debris, which provides sufficient data for constructing the SDTD dataset. Under long-exposure conditions in telescopic observations, debris often appears as a line source in the images. So, our simulation target is to generate realistic long-exposure line sources. We develop a pipeline that includes data pre-processing, line source debris video generation, and post processing. Fig 2 and the following paragraphs provide a detailed description of this process.
Figure 2: Overview of the space debris simulation pipeline. Step 1: Data Collection. Step 2: ZTF Raw Data Filtering. Step 3: Simulation Tool for Space Debris. Various parameters, including random initial points, angle change, length change, point spread function (PSF), fracture interference, and displacement, are applied to simulate space debris. The process generates synthetic video sequences and shows the real-world debris for reference.
@@ 3.1.2 Data Pre-Processing
where $z_{1}$ and $z_{2}$ represent the lower and upper bounds of the pixel intensity range, respectively. The hyperparameter $k$ is assigned a value of 2.5. Then, we perform a linear stretch on the grayscale values of the image, mapping the range $[z_{1},z_{2}]$ to [0,255][0,255][ 0 , 255 ]:
@@
$I_{\text{final}}(x,y)=\begin{cases}\max(I_{\text{scale}}(x,y),0),&I_{\text{scale}}(x,y)<0\\ \min(I_{\text{scale}}(x,y),255),&I_{\text{scale}}(x,y)>0\end{cases}$ (3)
@@ 3.2.2 Post Processing for Realistic
Since the real line source debris conforms to the Point Spread Function (PSF) characteristics [31], we perform point diffusion function processing on the preliminary line source debris. Among them, PSF is expressed as follows:
@@ 3.3.1 Dataset Statistics
After the simulation, we obtain 18,040 videos with a total of 62,562 frames. As shown in Fig. 12(a), to evaluate model performance in complex environments, we select 1,000 images from the 16,040 dataset to create two test sets: the debris test set and the dense debris test set. These two test sets allow for a comprehensive evaluation of model robustness under different conditions, including dense debris.
@@ 4.1 Overview
The entire pipeline of our Space Debris Tracking Network (SDT-Net) method is shown in Fig. 4, which is a CenterTrack-style [21] tracking method. Due to the debris long-exposure as linear, we treat the debris tracking as tracking for linear segments. For each object, we follow the tracking-by-detection idea. Our model detects two endpoints of each line and predicts velocity for cross-frame object association to generate trajectory.
@@ 4.3 Region-of-Interest Feature Enhancement
where $W,H$ are image dimensions, $M_{ij}$ is the ground truth, and $\hat{M}_{ij}$ is the predicted probability. The obtained $F_{en}$ is then passed through two modules: a Debris Detection module and a Debris Offset module.
@@ 5 Experiments
Table 2: Comparison of trackers on Debris Test and Dense Debris Test. The table compares the performance of various tracking algorithms in two distinct scenarios. The best and second--best figures are in color and $n$ represents the number of debris.
[…]
@@ 5.1 Experimental Setup
All experiments are performed using the mmdetection framework [39] with Pytorch [40]. The SDT-Net model is trained on 8 NVIDIA 4090 GPUs. The entire model is trained for 60 epochs, and the initial learning rate is set to 3e-3, which is decayed by 0.1 at the 20 epoch. The association radius $r$ to 200. Additionally, We adopt a downsampling strategy to generate paired images for training. Specifically, we first crop image patches of size 1524$\times$1524 from the original high-resolution images as ground truth. We employ the resolution of the input images is set to 1524$\times$1524 and the batch size is set to 2. Besides, the value of $\lambda_{\text{seg}}$ and $\lambda_{\text{emb}}$ are set to 1.0, $\lambda_{\text{hm}}$ is set to 10, and $\lambda_{\text{off}}$ is set to 0.1,
@@ 5.2 Evaluation Metrics
For evaluating the performance, we use standard MOT [41] evaluation metrics. Specifically, we adopt commonly used metrics in the field, including Multi-Object Tracking Accuracy (MOTA), identity F1 score (IDF1), Identity Switches (IDs), and Higher Order Tracking Accuracy (HOTA) [42]. These metrics provide a comprehensive assessment of tracking performance. Notably, HOTA includes Detection Accuracy (DetA) and Association Accuracy (AssA), where DetA measures the accuracy of detecting objects regardless of their identities, and AssA evaluates how well identities are consistently maintained across frames.
@@ 5.3 Comparison with Existing MOT Methods
Figure 5: Visualization example of the tracking results of the SDTD model on the test set. The green bounding box represents the GT, and the red bounding box represents the model prediction result. When the Intersection over Union (IoU) between the predicted box and GT is 0, the presence of only a green box indicates a False Negative, while the presence of only a red box indicates a False Positive.
@@ 5.4 Ablation Study
LSE RoI-FE Offset Module Debris Test ( $n\boldsymbol{\leq}2$) Dense Debris Test ($1\boldsymbol{\leq}n\boldsymbol{\leq}5$)
@@
Figure 6: Visualization of sample tracking results of real-world observations. Left: real-world debris observation. Right: model predicts tracking results, represented by predicted endpoints and connecting segments forming the trajectory (superimposed on a blank background). Different colors correspond to different debris IDs.
@@ 5.5 Qualitative Results
To further verify the effectiveness of SDT-Net, we evaluate the model with real-world debris data. Due to the scarcity of real-world debris observation data, we finally select 36 video sequences containing debris (a total of 2,228 frames) and invited astronomy experts to annotate the real-world observation data. As shown in Table 4, our method (SDT-Net) achieves a MOTA score exceeding 73%, demonstrating strong real-world performance. Fig 6 further illustrates the effectiveness of our approach. The left side of the figure presents raw debris observations, while the right side visualizes the predicted tracking results. Notably, the right side does not represent ground truth but rather our model’s predictions. For clarity, we display each tracked object using a line connecting its two key points instead of conventional bounding boxes. In practical applications, these tracked trajectories are overlaid on blank background images to facilitate interpretation. Each color in the visualization corresponds to a distinct debris ID, reflecting the model’s ability to distinguish and track multiple objects.
Table 4: Evaluation of SDT-Net on the Antarctic Station real-world data.
[…]
Deepsort [17] 48.7 48.9 51.8
Motdt [35] 47.4 45.2 47.6
CTracker [36] 51.9 45.8 54.5
Sort+YoloX [37] 57.4 57.4 58.9
CenterTrack [21] 64.7 58.3 68.7
ByteTrack [23] 66.8 62.4 70.1
OCSORT [38] 69.3 66.8 72.2
[…]